OpenStudy (anonymous):

Find the exact value of the expression: 2sin(17pi/2)-cos(2*13pi/3)+tan^2(pi/3)

OpenStudy (anonymous):

\[2\sin (\frac{ 17\pi }{ 2 })-\cos(2\frac{ 13\pi }{3 })+\tan^2(\frac{ \pi }{3 }) \]

OpenStudy (anonymous):

how do you do the math thing like the second post

OpenStudy (anonymous):

click where it says equation and you can type an equation

OpenStudy (anonymous):

lol wowowowowowoowwwwwwww. thanks bro.

OpenStudy (anonymous):

no problem